OpenStudy (anonymous):

3. Find the 7th term of the geometric sequence where t1 = 6 and r = 4.

OpenStudy (anonymous):

Well, the formula to find a term within a geometric sequence is: \[t_{n} = t_{1}(r)^{n-1}\] So, simply plug in your numbers: \[t_{n} = 6(4)^{7-1}\] \[t_{n} = 6(4)^{6}\] \[t_{n} = 6(4096)\] \[t_{n} = 24576\] Therefore, the 7th term is 24576